OpenStudy (anonymous):

What is the quantity of x in the equation x squared-13x+30=0

OpenStudy (anonymous):

You mean: \[x^{2}-13x+30\] right?

OpenStudy (raden):

use the formula : x1 + x2 = -b/a in this case, a = 1 and b = -13

OpenStudy (anonymous):

If so then: Factor it to get (x-10)(x-3) Then x=10 and 3.